OpenStudy (anonymous):

Solve B = –4(c – 2d) for d.

OpenStudy (texaschic101):

B = -4(c - 2d) --distribute the -4 through the parenthesis) B = -4c + 8d (add 4c to both sides) B + 4c = -4c+ 4c + 8d (combine like terms) B + 4c = 8d (now divide by 8) (B + 4c)/8 = d <-- your answer